### v1

# Towards Diverse and Efficient Audio Captioning via Diffusion Models

Abstract
We introduce Diffusion-based Audio Captioning (DAC), a non-autoregressive diffusion model tailored for diverse and efficient audio captioning. Although existing captioning models relying on language backbones have achieved remarkable success in various captioning tasks, their insufficient performance in terms of generation speed and diversity impede progress in audio understanding and multimedia applications. Our diffusion-based framework offers unique advantages stemming from its inherent stochasticity and holistic context modeling in captioning. Through rigorous evaluation, we demonstrate that DAC not only achieves SOTA performance levels compared to existing benchmarks in the caption quality, but also significantly outperforms them in terms of generation speed and diversity. The success of DAC illustrates that text generation can also be seamlessly integrated with audio and visual generation tasks using a diffusion backbone, paving the way for a unified, audio-related generative model across different modalities.  [^1] Project page at https://sites.google.com/view/diffusion-audio-captioning.
Index Terms:
audio captioning, diffusion model
I Introduction
Audio captioning involves detecting sound events and describing acoustic scenes using natural language. The community has witnessed remarkable achievements in audio captioning through Autoregressive (AR) models. Traditional encoder-decoder architectures [1, 2, 3, 4, 5] use audio encoders to extract audio features and leverage language decoders to generate coherent descriptions. More recently, Large Language Model (LLM)-based multimodal models [6, 7, 8] have emerged, driven by their superior captioning quality and diversity, thanks to a powerful language foundation. However, there are several minor yet non-negligible challenges associated with these models. Encoder-decoder-based models have a lower performance upper bound and can fall into the trap of generating monotonous and repetitive sentences due to their weaker decoders. LLM-based models are more powerful but require significantly more data and computational resources for training and have slower inference speeds due to the AR process.
In left-to-right generation tasks, such as text-to-audio or video-to-audio, diffusion models have emerged as a promising approach [9, 10, 11, 12], offering high-quality outputs and diversity. Furthermore, diffusion-based frameworks, due to the inherent advantages of Non-Autoregressive (NAR) models, excel at capturing the target-source dependency [13], resulting in a stronger connection between the input source media and the generated output. Although the NAR structure is typically thought unsuitable for generating inner-coherent contents like text, it facilitates faster generation speed due to parallel decoding and increased diversity due to stochastic noise sampling.
We consider extending diffusion-based frameworks to audio captioning tasks. In recent works, researchers have designed the pipeline for generating discrete text sequences based on various types of inputs as conditions. Specifically, Denoiser [14] facilitates diffusion models for discrete sequence generation by manipulating noises; LaDiC [15] revisits the advantages of diffusion models and highlights their strong competitiveness in image-to-text generation compared to AR models; Prefix-diffusion [16] proposes a lightweight diffusion model for diverse image captioning. A branch of works has demonstrated several key advantages of diffusion framework in text generation: 1) holistic context modeling, where models can capture more overall content rather than inner word relations; 2) parallel decoding, where tokens in the sequence are decoded in parallel and 3) diverse generation, which is brought from the diffusion models.
We propose DAC, a diffusion-based model specifically designed for diverse and efficient audio captioning. Building upon a wealth of research in image captioning and diffusion-based generation [17, 16], we develop DAC within the continuous text latent space. Initially, text descriptions are tokenized and transformed into discrete token embeddings, which are subsequently mapped into the continuous latent space in the form of word vectors using an embedding function. The audio content is converted into a Mel Spectrogram and encoded through a pre-trained audio encoder, followed by projection into the feature space via a lightweight projection module. The forward process adds noise to the text latent. In the backward process, the diffusion model predicts the noise added in each step, with the projected audio features conditioned as cross-attention. After denoising, text latent is decoded and converted back to discrete tokens through a language model head.
Figure 1: An overview of the proposed DAC framework. The audio is converted to the audio feature by a pre-trained audio encoder and then as the generation condition through cross attention. The diffusion backbone works on the continuous text latent space and then discrete into tokens.
Through evaluation, we demonstrate that DAC is not only competitive in terms of generation quality compared to SOTA baselines but also surpasses traditional AR methods in terms of generation diversity and speed. We also provide a further discussion of the commonly-used metrics in captioning tasks, revealing that DAC’s capabilities extend beyond these metrics. We incorporate extra semantic metrics such as CLAP and GPT4-eval to highlight DAC’s advantages. The effectiveness of diffusion-based captioning models paves the way towards a unified audio-related generation model, where we are possible to utilize a unified NAR structure for any-to-any conversion between different modalities, such as audio & text, and audio & image.
II Diffusion-based Audio Captioning (DAC)
II-A Diffusion Preliminaries
DAC is based on the Denoising Diffusion Probabilistic Model (DDPM) [18], where a forward process that repeatedly adds noise sampled from $N(0,I)$ to the input data:
$x_{t}=\sqrt{1-\beta_{t}}x_{t-1}+\sqrt{\beta_{t}}z_{t},$ (1)
where $x_{0}\sim q$ is the probability distribution to be learned, and then use an estimated function to estimate the undone process of each step in the forward diffusion process:
$p_{\theta}(x_{T})=N(x_{T}|0,I),$ (2)
$p_{\theta}(x_{t-1}|x_{t})=N(x_{t-1}|\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)),$ (3)
where the estimated function is parametrized by $\theta$, taking in two arguments $x_{t},t$ and outputting a vector $\mu_{\theta}(x_{t},t)$ and a matrix $\Sigma_{\theta(x_{t},t)}$. Maximum likelihood estimation with variational inference is commonly used to optimize the entire process. Traditional diffusion models use U-net [19] as the underlying architecture to predict noise, while modern models also employ transformer-based structures like DiT [20] and UViT [21]. The introduction of transformers allows for cross-attention as an effective way of incorporating conditions during generation.
II-B Discrete Text Diffusion Model
Diffusion models mainly work on the continuous latent space, while textual descriptions are discrete tokens. Works like D3PM [22] work on one-hot discrete vector input $x$, and use a categorical distribution $Cat(\cdot)$ and a transition matrix $[Qt]_{ij}=q(x_{t}=j|x_{t-1}=i)$ to model the probability:
$q(x_{t}|x_{t-1})=Cat(x_{t};p=x_{t-1}Q_{t}).$ (4)
A branch of other works like Diffusion-LM [17] and SSD-LM [23] mainly works on the continuous diffusion space. They have an extra embedding step in which discrete textual tokens $d={d_{i}}$ are embedded to continuous embedding $E(d)$:
$q_{\phi}(x_{0}|d)=N(E(d),\sigma_{0}I),$ (5)
and a final rounding step de-embed the continuous latent variable to discrete tokes:
$p_{\theta}(d|x_{0})=\prod_{i=1}^{n}p_{\theta}(d_{i}|x_{0}).$ (6)
DAC adopts continuous text diffusion models.
We follow LaDiC [15]’s design in the text decoder part, where we use BERT [24] to build the text decoder whose weights are initialized from the top 6 layers, and a following trainable language model (LM) head to refine the generated tokens.
In line with prior research in image captioning works [13, 15, 14, 16], our approach DAC primarily incorporates three types of losses: a Mean Squared Error (MSE) loss that measures the discrepancy between the original latent $x_{0}$ and the denoised latent $x^{\prime}_{0}$, a Cross Entropy (CE) loss that evaluates the alignment between the ground truth caption and the generated caption, and an auxiliary valid token loss that aids in constructing the output sequence. We have customized the loss function $\mathcal{L}$ to effectively balance the fitting of the denoised latent $x^{\prime}_{0}$ and the generation of the final textual description.
TABLE I: Evaluation of DAC versus SOTA baseline methods on the captioning QUALITY
Quality
Baseline Bleu_1 Bleu_4 METEOR Rouge CIDEr SPICE SPIDEr CLAP BERT GPT4-eval
ACT [25] 0.647 0.252 0.222 0.468 0.679 0.160 0.420 0.501 0.511 6.88
HTSAT-BART [3] 0.675 0.272 0.237 0.483 0.711 0.177 0.444 0.522 0.537 7.21
Prefix [4] 0.713 0.309 0.240 0.503 0.733 0.177 0.455 0.534 0.524 7.23
Pengi [5] 0.691 0.253 0.232 0.482 0.752 0.182 0.467 0.473 0.524 5.21
Audio-Flamingo [26] 0.449 0.079 0.191 0.344 0.266 0.124 0.195 0.461 0.484 5.76
Qwen-audio [7] 0.653 0.211 0.236 0.464 0.581 0.168 0.374 0.516 0.508 7.60
Qwen2-audio [8] 0.647 0.208 0.212 0.467 0.564 0.171 0.369 0.521 0.501 7.21
DAC-HTSAT 0.634 0.220 0.215 0.457 0.627 0.145 0.386 0.527 0.511 6.93
DAC-BEATs 0.672 0.221 0.226 0.460 0.611 0.154 0.382 0.553 0.508 7.11
DAC-BEATs-pt 0.711 0.295 0.251 0.492 0.655 0.172 0.474 0.548 0.536 7.37
II-C Audio Conditioning
DAC encodes audio information into the denoising process through cross-attention, where textual vectors serve as queries to retrieve hidden features from the audio latent space. We primarily leverage BEATs [27] as our audio feature encoder and use a projection module $\psi(\cdot)$ consisting of linear and layer normalization layers to map the audio features $A$ to the audio latent space. We also test other audio feature extractors such as Audio-Flamingo [26], HTS-AT [28], and Whipser [29]. Nevertheless, BEATs exhibits the best performance within our experiments.
The projected audio features $\psi(A)$ are introduced into the denoising process through cross-attention, which is integrated into each block of the denoising net. Mathematically, the cross-attention in DAC can be represented as:
$C_{t}^{c}=\text{Softmax}\left(\frac{Q_{x_{t}}\cdot K_{\psi(A)}}{\sqrt{d}}\right)\cdot V_{\psi(A)},$ (7)
where $Q_{x_{t}}$ is a projection of the continuous text embedding $x_{t}$, $K_{\psi(A)}$ and $V_{\psi(A)}$ are different projections of the projected audio features $\psi(A)$, and $d$ is the feature dimension of $K_{\psi(A)}$. The cross-attention mechanism allows text prompts to guide the generation process.
We also use classifier-free guidance to help improve the overall captioning quality [30]. During the inference time, the denoising process is performed both conditionally and unconditionally and then extrapolated according to a given weight $w$ called guidance scale:
$\tilde{\epsilon}_{\theta}=w\cdot\epsilon_{\theta}(x_{t},t,\psi(A))+(1-w)\cdot\epsilon_{\theta}(x_{t},t,\varnothing),$ (8)
where $\varnothing$ demotes the null audio feature.
III Experiments
III-A Datasets
We train our model and compare it with other baselines on the Audiocaps [31] dataset, which contains about 46k training instances, 3k validation instances and 3k evaluating instances. In the pre-training experiments, we also leverage Wavcaps [3] and Audioset [32] as extra training sets, which contain over 2 million training instances.
III-B Baselines
To the best of our knowledge, DAC is the first model utilizing the NAR framework for audio captioning task. We compare DAC with several SOTA AR audio captioning models, mainly falling into two categories: encoder-decoder-based models and LLM-based models. Specifically, ACT [33] trains a convolution-free transformer network to model the global information as well as capture temporal relationships in audio; HTSAT-BART [3] uses HTSAT [34] as the audio encoder and a pretrained BART [34] model as the language decoder; Pengi [5] frames all audio tasks as text-generation tasks through transfer learning; Prefix [4] leverages a mapping network to translate audio features to the continuous vectors so that language models can understand. At the billion-level parameter size, Audio-Flamingo [26] combines ClapCap [35] and Flamingo [36] for audio understanding and few-shot incontext learning and retrieval; Qwen-audio [7] and Qwen2-audio [8] are both built based on successful Qwen LLM foundations and show great audio understanding abilities.
For DAC models, we report two DAC models with different audio feature extractors: HTSAT (denoted as DAC-HTSAT), and BEATs (denoted as DAC-BEATs). We also evaluate a DAC-BEATs model that is firstly pre-trained on the Audioset and WavCaps datasets, and then fine-tuned on Audiocaps (denoted as DAC-BEATS-pt).
III-C Metrics
We evaluate several key aspects of the overall captioning capability of DAC compared with other baselines. Traditional captioning metrics focus on token-level matching from Natural Language Processing (NLP) level, including BLEU [37], ROGUEl [38], METEOR [39], CIDEr [40], SPICE [41], and SPIDEr [42]. Beyond these common metrics, we also use BERT [24] and CLAP [43] to compute the embedding similarity between text-text and text-audio as an extra evaluation of the overall semantic of the generated captions. Following the tendency in LLM evaluation, we further use GPT-4 to evaluate the generated caption quality. For the diversity, we mainly use two metrics: lexical-diversity (specifically, a measure of textual lexical diversity, MTLD [44]) and Distinct-N [45] (specifically, Distinct-1). MTLD measures the diversity of the textual lexicon, while Distinct-N measures the diversity of a sentence. We use Tokens Per Second (TPS) and Audios Per Second (APS) to measure the overall captioning speed of the baseline models.
III-D Experimental Settings
We train DAC on the Audiocaps dataset with an epoch of 80, a batch size of 128, and a learning rate of 1e-4, including 200 warming-up steps. We use Adam as the optimizer with betas = (0.9, 0.999). All training processes are conducted on 8 * Nvidia A100 cards. It takes about 50 GPU hours for the full training process. During the inference time, we set the inference skip step to 60 and the guidance scale to 2.5. We employ hyperparameter search methods to ensure the choice of parameters.
We test the inference speed on one single NVIDIA A100-SXM4-40GB card. Due to the different sizes of the checkpoints, we attempt to infer with the maximum batch size while loading all checkpoints with fp32. We run each experiment for 10 times, and report the average results.
IV Results and Analysis
IV-A Generation Quality
As shown in table I, we present the evaluation results of different captioning quality metrics in Audiocaps [^2] Some of the baseline scores are taken from their public releases. Comparing DAC with the other baselines, we demonstrate that although DAC is not the highest in the rankings among the baselines, it is competitive with the SOTA baselines in most metrics. The standard DAC model outperforms baselines like ACT and Qwen2-audio, and surpasses Pengi and HTSAT-BART in some of the quality metrics. The pre-trained version of DAC, DAC-BEATs-pt, significantly outperforms most of the baselines in all metrics. We would like to emphasize that since these metrics are built on sequences of tokens, AR models inherently possess advantages in these metrics compared to NAR models. In metrics like CLAP and BERT, we show that the DAC series of models greatly outperform the other baselines at a large scale, indicating that they have better captured the overall audio features and are more adept at generating holistic descriptions of the given audio.
Within the DAC groups, we illustrate two key findings: 1) the importance of the audio feature encoder in captioning, and 2) the significant impact of pre-training on enhancing captioning performance. Using HTSAT, a lightweight audio feature extractor trained for classification and sound event detection tasks, DAC-HTSAT performs competitively compared to most baselines. Furthermore, switching to BEATs as the audio encoder improves performance even more. We also observe that pre-training substantially boosts performance, consistent with the findings in the WavCaps paper. When comparing our results with those in WavCaps, we find that pre-training leads to a larger improvement in DAC (e.g., for the metric Bleu_1, 4.74% increase for HTSAT-BART and 5.80% for DAC-BEATs; for the metric SPIDEr, 9.23% increase for HTSAT-BART and 24.1% for DAC-BEATs). This can be attributed to the fact that AudioCaps is a relatively small dataset with a limited natural language corpus in the training set. AR models, especially those with a LLM as the language backbone, excel at understanding the intrinsic relationships within textual descriptions. In contrast, diffusion-based models require a larger pre-training corpus to effectively learn natural language.
TABLE II: Evaluation of DAC versus SOTA baseline methods on the captioning DIVERSITY and EFFICIENCY
Diversity Efficiency
Baseline Param MTLD Distinct-1 TPS APS
ACT [25] (beam=2) 384M 13.41 0.086 73.26 8.97
HTSAT-BART [3] 505M 13.12 0.078 67.92 5.78
Pengi [5] 550M 13.05 0.059 89.48 5.96
Audio-Flamingo [26] 2.2B 14.39 0.061 12.99 1.26
Qwen-audio [7] 7.7B 15.08 0.109 23.12 2.46
Qwen2-audio [8] 7.7B 14.67 0.089 25.28 2.43
DAC-HTSAT 573M 14.14 0.104 49.28 6.48
DAC-BEATs 627M 14.53 0.104 46.60 5.88
DAC-BEATs-pt 627M 14.45 0.100 47.52 6.64
IV-B Diversity and Efficiency
We present the diversity and efficiency evaluation results in table II, where we also report the model parameter size. We demonstrate that DAC models achieve significant diversity at both the lexical and sentence levels, while maintaining a relatively small model size and fast generation speed. In terms of diversity, only the Qwen-audio series surpass DAC, but with a much larger parameter size and a much slower inference speed. Audio-Flamingo outperforms ACT and HTSAT-BART with its pre-trained flamingo backbone, but lags behind the DAC models, particularly in terms of sentence diversity. We note that a stronger LLM backbone can indeed help increase generation diversity and quality, but this comes at the cost of greater computational resources and slower generation speed. Regarding generation efficiency, the fastest model, ACT, is almost eight times faster than the slowest Audio Flamingo with the same computational resources. Our DAC models are also significantly faster than AR models while maintaining competitive performance.
Beyond the quantitative results, we also observe two advantages that may not be evident from the table. Firstly, DAC is not sensitive to caption length due to their holistic denoising and decoding process on embeddings, as long as the length does not exceed the model’s token limit. This benefits both the generation speed and overall generation quality. For DAC, the generation token length limit is set to 40, while the average caption length for Audiocaps is around 10. In contrast, AR models experience a linear increase in generation time with sequence length. Longer sequences may also lead to distraction challenges in LLMs. Secondly, a smaller parameter size facilitates easier deployment on other devices or allows for larger batch sizes. Models with LLMs as backbones are challenging to deploy on most devices with billion-level parameter size.
Figure 2: Change of the metric scores on the Audiocaps dataset during fine-funing following a pre-training on the Audioset dataset. We demonstrate that achieving good convergence during Audioset pre-training does not necessarily translate to a strong initial performance on the AudioCaps test set, especially for metrics like BLEU and SPIDEr.
IV-C Further Discussion about the Metrics
An intriguing observation during the DAC fine-tuning process on Audiocaps is that achieving good convergence during the pre-training phase, e.g., pre-training on Audioset, does not guarantee a strong initial performance on the AudioCaps test set for certain metrics, as illustrated in fig. 2. Although pre-training on a significantly larger dataset yields much better results from the loss perspective (around 5 for Audiosat vs around 7 for AudioCaps), metrics such as Bleu, CIDEr, and SPIDEr are considerably lower than the baseline level if we directly evaluate the checkpoint on the Audiocaps dataset. In contrast, metrics like CLAP and BERT maintain consistently.
We thus consider metrics such as CLAP and BERT to be more equitable and convincing scores for evaluating captioning ability in our experiment. Traditional metrics like Bleu and CIDEr are based on token-level matching, implying that a better imitation of the ground truth strings results in a higher score. However, our findings suggest that a good caption does not necessarily have to replicate the ground truth verbatim. As demonstrated in fig. 2, there are instances where the generated captions are semantically similar to both the audio and the ground truth text, yet they receive low scores due to differences in caption style or grammar. Fine-tuning within the target distribution also does not necessarily enhance overall semantic similarity. On the other hand, metrics like CLAP and BERT measure the overall semantic similarity of the generated captions and the ground truth. On these two metrics, our model outperforms the other baselines.
V Conclusion
In this paper, we present Diffusion-based Audio Captioning (DAC), a state-of-the-art diffusion-based model for generating diverse and high-quality audio captioning. Leveraging the NAR architecture, DAC excels in providing a superior mix of diversity, efficiency, and quality in generating textual descriptions of audio pieces, outperforming existing AR models. Our model surpasses current benchmarks, especially in terms of generation diversity and processing speed, while maintaining a lightweight design. We hope that our findings will inspire further exploration of a diffusion-based comprehensive framework for multimodal content generation.

### v2

@@
 We introduce Diffusion-based Audio Captioning (DAC), a non-autoregressive diffusion model tailored for diverse and efficient audio captioning.
-Although existing captioning models relying on language backbones have achieved remarkable success in various captioning tasks, their insufficient performance in terms of generation speed and diversity impede progress in audio understanding and multimedia applications.
+Although existing captioning models relying on language backbones have achieved remarkable success in various captioning tasks, their insufficient performance in terms of generation speed and diversity impedes progress in audio understanding and multimedia applications.
 Our diffusion-based framework offers unique advantages stemming from its inherent stochasticity and holistic context modeling in captioning.
-Through rigorous evaluation, we demonstrate that DAC not only achieves SOTA performance levels compared to existing benchmarks in the caption quality, but also significantly outperforms them in terms of generation speed and diversity.
-The success of DAC illustrates that text generation can also be seamlessly integrated with audio and visual generation tasks using a diffusion backbone, paving the way for a unified, audio-related generative model across different modalities.
-[^1] Project page at https://sites.google.com/view/diffusion-audio-captioning.
-Index Terms:
+Through rigorous evaluation, we demonstrate that DAC not only achieves superior performance levels compared to existing benchmarks in the caption quality, but also significantly outperforms them in terms of generation speed and diversity.
+keywords:
 audio captioning, diffusion model
-I Introduction
+**footnotetext: Equal contribution\$$\dagger$\$$\dagger$\$footnotetext: Corresponding authorProject: https://sites.google.com/view/diffusion-audio-captioning
+1 Introduction
 Audio captioning involves detecting sound events and describing acoustic scenes using natural language.
@@
 Traditional encoder-decoder architectures [1, 2, 3, 4, 5] use audio encoders to extract audio features and leverage language decoders to generate coherent descriptions.
-More recently, Large Language Model (LLM)-based multimodal models [6, 7, 8] have emerged, driven by their superior captioning quality and diversity, thanks to a powerful language foundation.
+More recently, Large Language Model (LLM)-based multimodal models [6, 7] have emerged, driven by their superior captioning quality and diversity, thanks to a powerful language foundation.
 However, there are several minor yet non-negligible challenges associated with these models.
@@
 LLM-based models are more powerful but require significantly more data and computational resources for training and have slower inference speeds due to the AR process.
-In left-to-right generation tasks, such as text-to-audio or video-to-audio, diffusion models have emerged as a promising approach [9, 10, 11, 12], offering high-quality outputs and diversity.
-Furthermore, diffusion-based frameworks, due to the inherent advantages of Non-Autoregressive (NAR) models, excel at capturing the target-source dependency [13], resulting in a stronger connection between the input source media and the generated output.
+They may also suffer from hallucination problems [8].
+In left-to-right generation tasks, such as text-to-audio or video-to-audio, diffusion models have emerged as a promising approach [9, 10], offering high-quality outputs and diversity.
+Furthermore, diffusion-based frameworks, due to the inherent advantages of Non-Autoregressive (NAR) models, excel at capturing the target-source dependency [11, 12, 13], resulting in a stronger connection between the input source media and the generated output.
 Although the NAR structure is typically thought unsuitable for generating inner-coherent contents like text, it facilitates faster generation speed due to parallel decoding and increased diversity due to stochastic noise sampling.
-We consider extending diffusion-based frameworks to audio captioning tasks.
+A recent concurrent work DAC-RLD [14] leverages diffusion-based structure plus an AR decoder from Bart [15].
+The output decoding still follows the paradigm of the AR language model.
+As the fusion of AR and NAR, the generation is also affected and controlled by the AR decoder.
+We aim to extend diffusion-based NAR frameworks to audio captioning.
 In recent works, researchers have designed the pipeline for generating discrete text sequences based on various types of inputs as conditions.
-Specifically, Denoiser [14] facilitates diffusion models for discrete sequence generation by manipulating noises; LaDiC [15] revisits the advantages of diffusion models and highlights their strong competitiveness in image-to-text generation compared to AR models; Prefix-diffusion [16] proposes a lightweight diffusion model for diverse image captioning.
+Specifically, Denoiser [16] facilitates diffusion models for discrete sequence generation by manipulating noises; LaDiC [17] revisits the advantages of diffusion models and highlights their competitiveness in image-to-text generation compared to AR models; Prefix-diffusion [18] proposes a lightweight diffusion model for diverse image captioning.
 A branch of works has demonstrated several key advantages of diffusion framework in text generation: 1) holistic context modeling, where models can capture more overall content rather than inner word relations; 2) parallel decoding, where tokens in the sequence are decoded in parallel and 3) diverse generation, which is brought from the diffusion models.
-We propose DAC, a diffusion-based model specifically designed for diverse and efficient audio captioning.
-Building upon a wealth of research in image captioning and diffusion-based generation [17, 16], we develop DAC within the continuous text latent space.
-Initially, text descriptions are tokenized and transformed into discrete token embeddings, which are subsequently mapped into the continuous latent space in the form of word vectors using an embedding function.
-The audio content is converted into a Mel Spectrogram and encoded through a pre-trained audio encoder, followed by projection into the feature space via a lightweight projection module.
-The forward process adds noise to the text latent.
-In the backward process, the diffusion model predicts the noise added in each step, with the projected audio features conditioned as cross-attention.
-After denoising, text latent is decoded and converted back to discrete tokens through a language model head.
+We propose Diffusion-based Audio Captioning (DAC), a diffusion-based model for efficient audio captioning.
+Building on research in image captioning and diffusion-based generation [19, 18], DAC is a pure NAR model and operates in the continuous text latent space.
+Text descriptions are tokenized, embedded, and mapped into continuous word vectors.
+Audio is converted to a Mel Spectrogram, encoded by a pre-trained audio encoder, and projected into feature space.
+The forward process adds noise to the text latent, while the backward process uses the diffusion model to predict noise at each step, conditioned on projected audio features via cross-attention.
+After further post transition, the text latent is decoded into discrete tokens through a mapping model in a parallel way.
 Figure 1: An overview of the proposed DAC framework.
-The audio is converted to the audio feature by a pre-trained audio encoder and then as the generation condition through cross attention.
+The audio is converted to the audio feature and then as the generation condition through cross attention.
 The diffusion backbone works on the continuous text latent space and then discrete into tokens.
-Through evaluation, we demonstrate that DAC is not only competitive in terms of generation quality compared to SOTA baselines but also surpasses traditional AR methods in terms of generation diversity and speed.
+Through evaluation, we demonstrate that DAC is not only competitive in terms of generation quality compared to SOTA baselines but also surpasses several AR methods in terms of generation diversity and speed.
 We also provide a further discussion of the commonly-used metrics in captioning tasks, revealing that DAC’s capabilities extend beyond these metrics.
-We incorporate extra semantic metrics such as CLAP and GPT4-eval to highlight DAC’s advantages.
-The effectiveness of diffusion-based captioning models paves the way towards a unified audio-related generation model, where we are possible to utilize a unified NAR structure for any-to-any conversion between different modalities, such as audio & text, and audio & image.
-II Diffusion-based Audio Captioning (DAC)
-II-A Diffusion Preliminaries
+We incorporate extra semantic metrics such as CLAP [20], BERT [21], and GPT4-eval to highlight DAC’s advantages.
+2 Diffusion-based Audio Captioning
+2.1 Diffusion Preliminaries
 DAC is based on the Denoising Diffusion Probabilistic Model (DDPM) [22], where a forward process that repeatedly adds noise sampled from $N(0,I)$ to the input data:
@@
 Maximum likelihood estimation with variational inference is commonly used to optimize the entire process.
-Traditional diffusion models use U-net [19] as the underlying architecture to predict noise, while modern models also employ transformer-based structures like DiT [20] and UViT [21].
+Traditional diffusion models use U-Net as the underlying architecture to predict noise, while modern models also employ transformer-based structures like DiT [23] and UViT [24].
 The introduction of transformers allows for cross-attention as an effective way of incorporating conditions during generation.
-II-B Discrete Text Diffusion Model
+2.2 Discrete Text Diffusion Model
 Diffusion models mainly work on the continuous latent space, while textual descriptions are discrete tokens.
-Works like D3PM [22] work on one-hot discrete vector input $x$, and use a categorical distribution $Cat(\cdot)$ and a transition matrix $[Qt]_{ij}=q(x_{t}=j|x_{t-1}=i)$ to model the probability:
-$q(x_{t}|x_{t-1})=Cat(x_{t};p=x_{t-1}Q_{t}).$ (4)
-A branch of other works like Diffusion-LM [17] and SSD-LM [23] mainly works on the continuous diffusion space.
-They have an extra embedding step in which discrete textual tokens $d={d_{i}}$ are embedded to continuous embedding $E(d)$:
+Works like D3PM [25] work on one-hot discrete vector input, and use a categorical distribution and a transition matrix to model the transition probability.
+We adopt another branch of works like Diffusion-LM [19] and SSD-LM [26].
+The proposed framework DAC works on the continuous diffusion space.
+DAC has an extra embedding step in which discrete textual tokens $d={d_{i}}$ are embedded to continuous embedding $E(d)$:
 $q_{\phi}(x_{0}|d)=N(E(d),\sigma_{0}I),$ (4)
-and a final rounding step de-embed the continuous latent variable to discrete tokes:
+and a final rounding step de-embed the continuous latent variable to discrete tokens:
 $p_{\theta}(d|x_{0})=\prod_{i=1}^{n}p_{\theta}(d_{i}|x_{0}).$ (5)
-DAC adopts continuous text diffusion models.
-We follow LaDiC [15]’s design in the text decoder part, where we use BERT [24] to build the text decoder whose weights are initialized from the top 6 layers, and a following trainable language model (LM) head to refine the generated tokens.
-In line with prior research in image captioning works [13, 15, 14, 16], our approach DAC primarily incorporates three types of losses: a Mean Squared Error (MSE) loss that measures the discrepancy between the original latent $x_{0}$ and the denoised latent $x^{\prime}_{0}$, a Cross Entropy (CE) loss that evaluates the alignment between the ground truth caption and the generated caption, and an auxiliary valid token loss that aids in constructing the output sequence.
+DAC uses BERT [^2] https://huggingface.co/google-bert/bert-base-uncased [21] as the text encoder.
+The text decoder consists of two components, as shown in Fig.
+1: an embedding transition module that maps latent representations to the textual space, and a trainable language model (LM) head that converts embeddings into fluent text.
+For the transition module, we adopt the design from LaDiC [17], using BERT to construct the module, with weights initialized from the top layers.
+The LM head primarily consists of BERT PreTraining Heads, along with additional linear layers, which serve as the generator for discrete tokens.
+The diffusion module is implemented in two versions: a UViT-based and a DiT-based framework.
+Similar to image captioning works [11, 17, 16, 18], DAC primarily incorporates three types of losses: a Mean Squared Error loss that measures the discrepancy between the original latent $x_{0}$ and the denoised latent $x^{\prime}_{0}$, a Cross Entropy loss that evaluates the alignment between the ground truth caption and the generated caption, and an auxiliary valid token loss that aids in constructing the output sequence.
 We have customized the loss function $\mathcal{L}$ to effectively balance the fitting of the denoised latent $x^{\prime}_{0}$ and the generation of the final textual description.
-TABLE I: Evaluation of DAC versus SOTA baseline methods on the captioning QUALITY
+Table 1: Evaluation of DAC versus SOTA baseline methods on the captioning QUALITY (higher is better)
 Quality
@@
 Pengi [5] 0.691 0.253 0.232 0.482 0.752 0.182 0.467 0.473 0.524 5.21
+DAC-RLD [14] 0.671 0.279 0.255 0.497 0.755 0.187 0.471 0.531 0.521 7.12
 Audio-Flamingo [28] 0.449 0.079 0.191 0.344 0.266 0.124 0.195 0.461 0.484 5.76
 Qwen-audio [6] 0.653 0.211 0.236 0.464 0.581 0.168 0.374 0.516 0.508 7.60
-Qwen2-audio [8] 0.647 0.208 0.212 0.467 0.564 0.171 0.369 0.521 0.501 7.21
-DAC-HTSAT 0.634 0.220 0.215 0.457 0.627 0.145 0.386 0.527 0.511 6.93
-DAC-BEATs 0.672 0.221 0.226 0.460 0.611 0.154 0.382 0.553 0.508 7.11
-DAC-BEATs-pt 0.711 0.295 0.251 0.492 0.655 0.172 0.474 0.548 0.536 7.37
-II-C Audio Conditioning
+Qwen2-audio [7] 0.647 0.208 0.212 0.467 0.564 0.171 0.369 0.521 0.538 7.36
+DAC (HTSAT, UViT) 0.634 0.220 0.215 0.457 0.627 0.145 0.386 0.527 0.511 6.93
+DAC (HTSAT, DiT) 0.638 0.235 0.223 0.449 0.631 0.144 0.392 0.522 0.506 6.72
+DAC (BEATs, UViT) 0.672 0.221 0.226 0.460 0.611 0.154 0.382 0.553 0.508 7.11
+DAC (BEATs, UViT, pt) 0.711 0.295 0.251 0.492 0.655 0.172 0.414 0.548 0.536 7.37
+DAC (BEATs, DiT) 0.674 0.233 0.226 0.455 0.635 0.148 0.404 0.546 0.511 6.92
+DAC (BEATs, DiT, pt) 0.713 0.298 0.253 0.488 0.674 0.178 0.421 0.549 0.529 7.21
+2.3 Audio Conditioning
 DAC encodes audio information into the denoising process through cross-attention, where textual vectors serve as queries to retrieve hidden features from the audio latent space.
-We primarily leverage BEATs [27] as our audio feature encoder and use a projection module $\psi(\cdot)$ consisting of linear and layer normalization layers to map the audio features $A$ to the audio latent space.
-We also test other audio feature extractors such as Audio-Flamingo [26], HTS-AT [28], and Whipser [29].
-Nevertheless, BEATs exhibits the best performance within our experiments.
+We primarily leverage BEATs [29] and HTSAT [30] as our audio feature encoder and use a projection module $\psi(\cdot)$ consisting of linear and layer normalization layers to map the audio features $A$ to the audio latent space.
 The projected audio features $\psi(A)$ are introduced into the denoising process through cross-attention, which is integrated into each block of the denoising net.
@@
 where $\varnothing$ demotes the null audio feature.
-III Experiments
-III-A Datasets
-We train our model and compare it with other baselines on the Audiocaps [31] dataset, which contains about 46k training instances, 3k validation instances and 3k evaluating instances.
+3 Experiments
+3.1 Datasets
+We train our model and compare it with other baselines on the AudioCaps [32] dataset, which contains about 46k training instances, 447 validation instances and 957 evaluating instances.
 In the pre-training experiments, we also leverage Wavcaps [3] and Audioset [33] as extra training sets, which contain over 2 million training instances.
-III-B Baselines
-To the best of our knowledge, DAC is the first model utilizing the NAR framework for audio captioning task.
-We compare DAC with several SOTA AR audio captioning models, mainly falling into two categories: encoder-decoder-based models and LLM-based models.
-Specifically, ACT [33] trains a convolution-free transformer network to model the global information as well as capture temporal relationships in audio; HTSAT-BART [3] uses HTSAT [34] as the audio encoder and a pretrained BART [34] model as the language decoder; Pengi [5] frames all audio tasks as text-generation tasks through transfer learning; Prefix [4] leverages a mapping network to translate audio features to the continuous vectors so that language models can understand.
-At the billion-level parameter size, Audio-Flamingo [26] combines ClapCap [35] and Flamingo [36] for audio understanding and few-shot incontext learning and retrieval; Qwen-audio [7] and Qwen2-audio [8] are both built based on successful Qwen LLM foundations and show great audio understanding abilities.
-For DAC models, we report two DAC models with different audio feature extractors: HTSAT (denoted as DAC-HTSAT), and BEATs (denoted as DAC-BEATs).
-We also evaluate a DAC-BEATs model that is firstly pre-trained on the Audioset and WavCaps datasets, and then fine-tuned on Audiocaps (denoted as DAC-BEATS-pt).
-III-C Metrics
-We evaluate several key aspects of the overall captioning capability of DAC compared with other baselines.
-Traditional captioning metrics focus on token-level matching from Natural Language Processing (NLP) level, including BLEU [37], ROGUEl [38], METEOR [39], CIDEr [40], SPICE [41], and SPIDEr [42].
-Beyond these common metrics, we also use BERT [24] and CLAP [43] to compute the embedding similarity between text-text and text-audio as an extra evaluation of the overall semantic of the generated captions.
-Following the tendency in LLM evaluation, we further use GPT-4 to evaluate the generated caption quality.
+3.2 Baselines
+We compare DAC with several open-source SOTA AR audio captioning models.
+Specifically, we compare encoder-decoder-based models like ACT [27], HTSAT-BART [3], Pengi [5], and Prefix [4].
+At the billion-level parameter size, we compare LLM-based models like Audio-Flamingo [28], Qwen-audio [6] and Qwen2-audio [7].
+We also compare diffusion-based concurrent work DAC-RLD [14].
+For DAC models, we report two DAC models with different audio feature extractors: HTSAT (HTSAT-Audioset), and BEATs (BEATs-ft).
+We also evaluate a DAC model that is firstly pre-trained on the Audioset and Wavcaps datasets, and then fine-tuned on AudioCaps.
+3.3 Metrics
+We evaluate several key aspects of the captioning capability of DAC compared with other baselines.
+Traditional metrics focus on token-level matching, including BLEU, ROGUEl, METEOR, CIDEr, SPICE, and SPIDEr.
+Beyond these common metrics, we also use BERT (bert-base-uncased) [21] and CLAP [^3] https://huggingface.co/lukewys/laion_clap/blob/main/630k-best.pt [20] to compute the embedding similarity between text-text and text-audio as an extra evaluation of the overall semantic of the generated captions.
+Following the tendency in LLM evaluation, we further use GPT-4 (gpt-4-0613) to evaluate the generated caption quality.
 For the diversity, we mainly use two metrics: lexical-diversity (specifically, a measure of textual lexical diversity, MTLD [34]) and Distinct-N [35] (specifically, Distinct-1).
 MTLD measures the diversity of the textual lexicon, while Distinct-N measures the diversity of a sentence.
-We use Tokens Per Second (TPS) and Audios Per Second (APS) to measure the overall captioning speed of the baseline models.
-III-D Experimental Settings
-We train DAC on the Audiocaps dataset with an epoch of 80, a batch size of 128, and a learning rate of 1e-4, including 200 warming-up steps.
-We use Adam as the optimizer with betas = (0.9, 0.999).
-All training processes are conducted on 8 * Nvidia A100 cards.
-It takes about 50 GPU hours for the full training process.
-During the inference time, we set the inference skip step to 60 and the guidance scale to 2.5.
-We employ hyperparameter search methods to ensure the choice of parameters.
-We test the inference speed on one single NVIDIA A100-SXM4-40GB card.
-Due to the different sizes of the checkpoints, we attempt to infer with the maximum batch size while loading all checkpoints with fp32.
-We run each experiment for 10 times, and report the average results.
-IV Results and Analysis
-IV-A Generation Quality
-As shown in table I, we present the evaluation results of different captioning quality metrics in Audiocaps [^2] Some of the baseline scores are taken from their public releases.
+We use Tokens Per Second (TPS) and Audios Per Second (APS) to measure the overall captioning speed of the models.
+3.4 Experimental Settings
+We train DAC on the AudioCaps for 80 epochs with a batch size of 128 and a learning rate of 1e-4, including 200 warm-up steps.
+The base model is built on 2D UViT with 6 layers for each encoder and decoder, and the DiT model is based on a 12-layer Transformer with 768 hidden sizes and 12 attention heads.
+We use Adam as the optimizer and train on 8×NVIDIA A100 GPUs.
+During inference, we set the skip step to 60 and the guidance scale to 2.5, optimizing parameters via hyperparameter search.
+For the pre-trained version, We pre-train DAC on Audioset and Wavcaps for 60 epochs, and then finetune on AudioCaps for 20 epochs.
+Inference speed is tested on a single NVIDIA A100-SXM4-40GB.
+We use the largest possible batch size while loading checkpoints in fp32.
+Each experiment runs 10 times, and we report the average results.
+4 Results and Analysis
+4.1 Generation Quality
+As shown in table 1, we present the evaluation results of different captioning quality metrics in AudioCaps [^4] Baseline scores are taken from original papers or implemented and benchmarked using the official public repositories..
 Comparing DAC with the other baselines, we demonstrate that although DAC is not the highest in the rankings among the baselines, it is competitive with the SOTA baselines in most metrics.
-The standard DAC model outperforms baselines like ACT and Qwen2-audio, and surpasses Pengi and HTSAT-BART in some of the quality metrics.
-The pre-trained version of DAC, DAC-BEATs-pt, significantly outperforms most of the baselines in all metrics.
-We would like to emphasize that since these metrics are built on sequences of tokens, AR models inherently possess advantages in these metrics compared to NAR models.
-In metrics like CLAP and BERT, we show that the DAC series of models greatly outperform the other baselines at a large scale, indicating that they have better captured the overall audio features and are more adept at generating holistic descriptions of the given audio.
-Within the DAC groups, we illustrate two key findings: 1) the importance of the audio feature encoder in captioning, and 2) the significant impact of pre-training on enhancing captioning performance.
-Using HTSAT, a lightweight audio feature extractor trained for classification and sound event detection tasks, DAC-HTSAT performs competitively compared to most baselines.
-Furthermore, switching to BEATs as the audio encoder improves performance even more.
-We also observe that pre-training substantially boosts performance, consistent with the findings in the WavCaps paper.
-When comparing our results with those in WavCaps, we find that pre-training leads to a larger improvement in DAC (e.g., for the metric Bleu_1, 4.74% increase for HTSAT-BART and 5.80% for DAC-BEATs; for the metric SPIDEr, 9.23% increase for HTSAT-BART and 24.1% for DAC-BEATs).
-This can be attributed to the fact that AudioCaps is a relatively small dataset with a limited natural language corpus in the training set.
-AR models, especially those with a LLM as the language backbone, excel at understanding the intrinsic relationships within textual descriptions.
-In contrast, diffusion-based models require a larger pre-training corpus to effectively learn natural language.
-TABLE II: Evaluation of DAC versus SOTA baseline methods on the captioning DIVERSITY and EFFICIENCY
+The standard DAC model outperforms baselines like ACT and Qwen2-audio and surpasses Pengi and HTSAT-BART in some of the quality metrics.
+The pre-trained version of DAC (denoted as pt) outperforms most of the baselines on average across the metrics (We demonstrate that no single model excels in every metric).
+We would also like to emphasize that since these metrics are built on sequences of tokens, AR models inherently possess advantages in these metrics compared to NAR models.
+In metrics like CLAP and BERT, the DAC series of models greatly outperform the other baselines at a large scale, indicating that they have better captured the overall audio features and are more adept at generating holistic descriptions of the given audio.
+Within the DAC groups, we highlight two key findings: 1) the importance of the audio feature encoder for captioning and 2) the significant impact of pre-training on performance.
+Using HTSAT, a lightweight audio extractor trained for classification and sound event detection, DAC (UViT, HTSAT) performs competitively with most baselines.
+Switching to BEATs further improves performance.
+We also observe that pre-training boosts performance, as shown in the Wavcaps paper.
+Comparing our results with Wavcaps, we find that pre-training results in a larger improvement in DAC (e.g., BLEU_1 increases by 4.74% for HTSAT-BART and 5.80% for DAC, while SPIDEr increases by 9.23% for HTSAT-BART and 24.1% for DAC).
+This can be attributed to AudioCaps’ relatively small dataset with a limited language corpus.
+AR models, especially those with LLMs, excel at understanding textual relationships, while diffusion-based models require a larger corpus for effective language learning.
+Table 2: Evaluation of DAC versus SOTA baseline methods on the captioning DIVERSITY and EFFICIENCY (higher is better)
 Diversity Efficiency
@@
 Pengi [5] 550M 13.05 0.059 89.48 5.96
+DAC-RLD [14] 372M 14.08 0.089 63.42 6.97
 Audio-Flamingo [28] 2.2B 14.39 0.061 12.99 1.26
@@
 Qwen2-audio [7] 7.7B 14.67 0.089 25.28 2.43
-DAC-HTSAT 573M 14.14 0.104 49.28 6.48
-DAC-BEATs 627M 14.53 0.104 46.60 5.88
-DAC-BEATs-pt 627M 14.45 0.100 47.52 6.64
-IV-B Diversity and Efficiency
-We present the diversity and efficiency evaluation results in table II, where we also report the model parameter size.
+DAC (HTSAT, UViT) 573M 14.14 0.104 49.28 6.48
+DAC (HTSAT, DiT) 597M 14.12 0.098 67.32 8.29
+DAC (BEATs, UViT) 627M 14.53 0.104 46.60 5.88
+DAC (BEATs, UViT, pt) 627M 14.45 0.100 47.52 6.64
+DAC (BEATs, DiT) 651M 14.50 0.101 60.28 8.03
+DAC (BEATs, DiT, pt) 651M 14.53 0.102 60.23 8.12
+4.2 Diversity and Efficiency
+We present the diversity and efficiency evaluation results in table 2, where we also report the model parameter size.
 We demonstrate that DAC models achieve significant diversity at both the lexical and sentence levels, while maintaining a relatively small model size and fast generation speed.
 In terms of diversity, only the Qwen-audio series surpass DAC, but with a much larger parameter size and a much slower inference speed.
-Audio-Flamingo outperforms ACT and HTSAT-BART with its pre-trained flamingo backbone, but lags behind the DAC models, particularly in terms of sentence diversity.
-We note that a stronger LLM backbone can indeed help increase generation diversity and quality, but this comes at the cost of greater computational resources and slower generation speed.
+A stronger LLM backbone can indeed help increase generation diversity and quality, but this comes at the cost of greater computational resources and slower generation speed.
 Regarding generation efficiency, the fastest model, ACT, is almost eight times faster than the slowest Audio Flamingo with the same computational resources.
 Our DAC models are also significantly faster than AR models while maintaining competitive performance.
-Beyond the quantitative results, we also observe two advantages that may not be evident from the table.
+Compared to DAC-RLD, our model is competitive in generation speed but outperforms it in diversity.
+We attribute this to the DAC-RLD design, where the subsequent AR decoder from the BART module may mask the potential of diffusion in enhancing diversity.
+Beyond the quantitative results, we observe two advantages that may not be evident from the table.
 Firstly, DAC is not sensitive to caption length due to their holistic denoising and decoding process on embeddings, as long as the length does not exceed the model’s token limit.
 This benefits both the generation speed and overall generation quality.
-For DAC, the generation token length limit is set to 40, while the average caption length for Audiocaps is around 10.
+For DAC, the generation token length is limited to 40, while the average caption length for AudioCaps is around 10.
 In contrast, AR models experience a linear increase in generation time with sequence length.
 Longer sequences may also lead to distraction challenges in LLMs.
-Secondly, a smaller parameter size facilitates easier deployment on other devices or allows for larger batch sizes.
-Models with LLMs as backbones are challenging to deploy on most devices with billion-level parameter size.
-Figure 2: Change of the metric scores on the Audiocaps dataset during fine-funing following a pre-training on the Audioset dataset.
+Secondly, a smaller parameter size facilitates easier deployment on devices or allows for larger batch sizes.
+Figure 2: Change of the metric scores on the AudioCaps dataset during fine-funing following pre-training.
 We demonstrate that achieving good convergence during Audioset pre-training does not necessarily translate to a strong initial performance on the AudioCaps test set, especially for metrics like BLEU and SPIDEr.
-IV-C Further Discussion about the Metrics
+4.3 Further Discussion about the Metrics
 An intriguing observation during the DAC fine-tuning process on AudioCaps is that achieving good convergence during the pre-training phase, e.g., pre-training on Audioset, does not guarantee a strong initial performance on the AudioCaps test set for certain metrics, as illustrated in fig.
 2.
-Although pre-training on a significantly larger dataset yields much better results from the loss perspective (around 5 for Audiosat vs around 7 for AudioCaps), metrics such as Bleu, CIDEr, and SPIDEr are considerably lower than the baseline level if we directly evaluate the checkpoint on the Audiocaps dataset.
-In contrast, metrics like CLAP and BERT maintain consistently.
+Although pre-training on a significantly larger dataset yields much better results from the loss perspective (around 5 for Audioset vs around 7 for AudioCaps), metrics such as BLEU, CIDEr, and SPIDEr are considerably lower than the baseline level if we directly evaluate the checkpoint on the AudioCaps dataset.
+In contrast, metrics like CLAP and BERT are maintained consistently.
 We thus consider metrics such as CLAP and BERT to be more equitable and convincing scores for evaluating captioning ability in our experiment.
@@
 Fine-tuning within the target distribution also does not necessarily enhance overall semantic similarity.
-On the other hand, metrics like CLAP and BERT measure the overall semantic similarity of the generated captions and the ground truth.
+On the other hand, metrics like CLAP and BERT measure the overall semantic similarity.
 On these two metrics, our model outperforms the other baselines.
-V Conclusion
-In this paper, we present Diffusion-based Audio Captioning (DAC), a state-of-the-art diffusion-based model for generating diverse and high-quality audio captioning.
-Leveraging the NAR architecture, DAC excels in providing a superior mix of diversity, efficiency, and quality in generating textual descriptions of audio pieces, outperforming existing AR models.
+5 Conclusion
+We present DAC, a diffusion-based NAR model for generating diverse and high-quality audio captioning.
+Leveraging the NAR architecture, DAC excels in providing a superior mix of diversity, efficiency, and quality in generating textual descriptions of audio pieces, outperforming several existing AR models.
 Our model surpasses current benchmarks, especially in terms of generation diversity and processing speed, while maintaining a lightweight design.